For a particular product you have a choice of two designs. The first design will result in a non-recurring engineering cost (NRE cost) of $30,000 and a unit cost of $50, the second design will result in a non-recurring engineering cost (NRE cost) of $200,000 and a unit cost of $10.

How many units do you need to sell before the second design becomes more cost effective?
Answer #1

Answer:

Non recurring engineering cost is a one time cost
Let x be the no. of units need to sold before second design become more cost effective.
when x no. of units sold the cost of both designs are same
Then,
Cost of x units from first design = cost of x unit from second design

30000 + 50x = 200000 + 10x
40x = 170000
x = 170000 / 40
x = 4250

Now let us see the cost for both design when x+1 i.e 4251 units sold
cost of first design = 30000 + 50*4251
= 2,42,550‬

cost of second design = 200000 + 10*4251
= 2,42,510

Its clear from here that for x+1 i.e 4251 units second design is more cost effective

Now let use see the cost of both design when x-1 i.e 4249 units sold
cost of first design = 30000 + 50*4249
= 2,42,450

cost of second design = 200000 + 10* 4249
= 2,42,490

Its clear from here that for x-1 i.e 4249 units first design is more cost effective

There 4250 units need to be sold before the second design become more cost effective

Please upvote if you like the answer!